OpenStudy (anonymous):

find angle supplementary to θ. a)θ=125°16'27" b) 58.07° answer: a) 54°43'33" b)121.93° find exact radian measure of the angle 150° ans:(5π/6) find exact degree measure of the angle 2π/3 ans: 120° I just need some help on how to answer these last questions for this math section, I do not know how to get these answers! Please help! Thanks for any help! :D

OpenStudy (anonymous):

first conver it to decimal . 125.27 subtract from 180 180-125.27=54.73 convert it back. a) 54°43'33

OpenStudy (anonymous):

you need help with only 2pi/ 3?

terenzreignz (terenzreignz):

Seems to be off line

OpenStudy (anonymous):

hes right

OpenStudy (anonymous):

since pi radion =180 \[\Large 1 radion =\frac{180}{\pi}\] for 2pi/3 \[\Large =\frac{180}{\pi}*\frac{2\pi}{3}=120\]

OpenStudy (anonymous):

Sorry it was late when I posted this and since no one was answering for a while I just decided to go to sleep. As for your question, I didn't only need help with "2pi/ 3" there were few other problems, but I just want to know how to answer these questions.

OpenStudy (anonymous):

for your first response, when you said you get "125.27" where does the ".27" come from and what do you mean "convert it back"? Im not sure how to get those answers :O

OpenStudy (anonymous):

well to convert degree minutes and seconds to decimal Divide the Minutes with 60 and divide the seconds with 3600 and add them to degrees here 16 minutes so 16/60=0.26666666666666666666 here 27 sexconds 27/3600=0.0075 add them 125+0.2667+0.0075=125.2742 to convert back to degree minutes seconds take the decimal part here .2742 Multiply with 60 0.2742*60=16.452 here the whole numers are minutes take decimal part from above .452 and Multiply with 60 0.452*60=27.12 these are second so 125°16' 27"